OpenStudy (anonymous):

Suppose the diameter of a circle is 9. What is its circumference?

OpenStudy (anonymous):

If you could give the answer in terms of pi that'd be great.

sammixboo (sammixboo):

Well we must first find the radius of the circle. The radius is half of the diameter, so our radios os 4.5 r = 4.5 Formula to find the Circumference: \(2\pi*4.5\)

sammixboo (sammixboo):

\(C~=~2\pi*4.5\)

OpenStudy (anonymous):

circumference equals pi times the diameter.

sammixboo (sammixboo):

Do you know what to do next? And the circumference equals pi times the radius, not the diameter

OpenStudy (anonymous):

circumference is 28.26

OpenStudy (anonymous):

Thanks. Could you help me with a few more?

sammixboo (sammixboo):

Just multiply 4.5 by 2, so what is 4.5 x 2?

OpenStudy (anonymous):

9pi, I know.

sammixboo (sammixboo):

\(C~=~9\pi\) Or \(C~=~9*3.14\)

OpenStudy (anonymous):

Can you help me with a few more please?

sammixboo (sammixboo):

Sure I can try :)

sammixboo (sammixboo):

I might take a bit to answer here and there.. I am doing a test right now :(

OpenStudy (anonymous):

Suppose the circumference of a circle is 15π. What is its diameter?

OpenStudy (anonymous):

Those question marks are supposed to be pi. |:(

OpenStudy (anonymous):

I don't know, it's not multiple choice and gives practically no info on how to answer.

sammixboo (sammixboo):

C = 2pi*r 47.1 = 6.28*r Now we must find the raduis to find the diameter. To do so we must divide 6.28 on both sides of the equation. 47.1/6.28 = 6.28*r/6.28 47.1/6.28 = r What is 47.1/6.28?

sammixboo (sammixboo):

15pi = 47.1, sorry I completely skipped a step >_< And 2*pi = 6.28

sammixboo (sammixboo):

After you find out what is 47.1/6.28 you multiply the answer by 2, since the radius is half of the diameter, and that is your answer

OpenStudy (anonymous):

Thanks, you've just helped me a lot. :D